• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Seonho Choi

    Log entry time 15:50:16 on February19,2002

    Entry number 83201

    keyword=Shift Summary (Feb. 19 day)

    08:00 Just finished 3B(b), moved the target from optics to LH2 no rad
    08:15 Asked MCC 25uA for 3B(c) setting, run 2678 started
    08:30 Bigbox magnet on going for 3B(d)
    09:00 3B(d) finished. Target moved to LH2 rad, Spetro rotated to 32.46, momentum changed to 2.142 GeV/c
    09:15 Asked 5uA and start run 2680
    11:00 Found out that the spectrometer angle is 0.5 degrees off from the plan. Fixed it.
    14:00 Veto HV turned on, VV12~2MHz at 5uA
    14:45 Run 2686 stop failed reboot TS2, kcoda and start from scratch
    14:55 Calo magnet off, veto HV off, start 3B(f) with run # 2687
    15:35 3B(f) finished and moving to optics target.
    15:40 Asked 25uA without raster and optics run 2688 started.
    15:50 Finished everything for kinematics 3B. Asked controlled acess to the hall.